OpenStudy (anonymous):

LIL BROS MATHS: I NEED HELP. Use algebra to determine the dimensions of a spare based popper box that will hold 250ml, yet has the smallest possible surface area. Use a graphics calculator and show all working out........anyone know how to do this properly? There are lots of answers.....and no google help...

OpenStudy (anonymous):

Can I assume that W = H = L

OpenStudy (anonymous):

dude....no its a square based prism.

OpenStudy (anonymous):

SA = 2 * b * sqrt(b^2+h^2/4) + b^2 V = 1/3 * b^2 * h 250 = 1/3 * b^2 * h b = 5 * sqrt(30) / sqrt(h) Substitute b into SA, calculate minimum of derivative in respect to h

OpenStudy (anonymous):

ok this is what we had so far... S = 2x^2 + 250/x after substituting for x, s==94.5 and x=4.... didnt seem right... going to try your method now....assuming SA is surface area and b= ? sprt=?

OpenStudy (anonymous):

sqrt is square root b is base length

OpenStudy (anonymous):

ok cool.... so then to find the smallest surface area, we substitute into b? (im not fantastic at maths) we have to make a graph and a table etc...

OpenStudy (anonymous):

you have b = 5 * sqrt(30) / sqrt(h) so you substitute each b in SA to that, find derivative, and then the minimum

OpenStudy (anonymous):

ahhh ok........going to work on it right now

OpenStudy (anonymous):

soo whats the answer.....whats the smallest surface area in your workin out